# A Primal-Dual-Assisted Penalty Approach to Bilevel Optimization with Coupled Constraints

### v1

@@ 4.1 Toy example
Table 2: Numerical results on the training outcome of our BLOCC with LV-HBA [72] and GAM [69]. The first row represents accuracy $\pm$ standard deviation, and the second row between brackets represents the running time until the upper-level objective’s update is smaller than 1e-5.
@@ 4.2 Hyperparameter optimization for SVM
We test the performance of our algorithm BLOCC with $\gamma=10$ when training a linear SVM model on the diabetes [20] and fourclass datasets [29]. The model is trained via the BLO formulation (1) with inequality CCs; see the details in Appendix E. We compared performance with two baselines, LV-HBA [72] and GAM [69].
@@ 4.3 Transportation network design problem
BLO is particularly relevant in transportation, where network planning must consider different time horizons and actors. Those problems have a large number of CCs, challenging the use of traditional BLO techniques. As a result, our final experiment considers a network design problem, where we act as an operator (whose utility is modeled as the upper-level objective) that considers the passengers’ behavior, modeled in the lower level. We considered three networks: two synthetic networks of 3 and 9 nodes, respectively, and a real-world network of 26 nodes in the city of Seville, Spain. Further details about the formulation and the experiment can be found in Appendix F.
In this experiment, we only compare our BLOCC with LV-HBA, as GAM involves calculating the inverse of the Hessian, which is not tractable in this setting. Table 3 summarizes the experiment results. We can see that LV-HBA failed to work efficiently, especially for large networks. This is mainly because the increased constraints render the projection step impracticable. Our BLOCC, in contrast, is much faster, and it successfully converges even with large real-world networks.
@@ B.1 Proof of Lemma 1
As $g(x,y)$ is $\alpha_{g}$-strongly convex in $y$ and $g^{c}(x,y)$ is convex in $y$, we know $g(x,y)+\langle\mu_{g}^{*}(x),g^{c}(x,y)\rangle$ is $\alpha_{g}$-strongly convex in $y$, where the modulus $\alpha_{g}$ is independent of $x$. Therefore, according to Theorem 2 in [37], the quadratic growth in statement 1 can be concluded.
@@ D.3 Proof of Theorem 4
with complexity $T={\cal O}(\log(\epsilon^{-1}))$. Here, $(\mu^{*},y^{*})=\arg\max_{\mu\in\mathbb{R}^{d_{c}}}\min_{y\in\mathbb{R}^{d_{y}}}L(\mu,y)$.
@@
Notice ${\cal O}(\alpha_{h_{2}}^{-1})<{\cal O}(1)$ and $P_{0}={\cal O}(\epsilon^{-1})$ as $\rho=\epsilon^{-1}$. In this way, there exist $T_{1}={\cal O}(\log(\epsilon^{-1}))$ such that for all $t>T_{1}$, $(1-c)^{t}P_{0}={\cal O}(1)$ and ${\cal O}(\alpha_{h_{2}}^{-1})\leq{\cal O}(1)$. Accordingly, we can achieve
@@
This is an iteration outcome using (50) in Lemma 12, (62), and the fact that $\eta_{1}\alpha_{h_{2}}/2={\cal O}(\frac{\alpha_{h_{2}}}{l_{h_{2},1}})={\cal O}(1)$. In this way, for another $T_{2}={\cal O}(\log(\epsilon^{-1}))$ steps, we have
@@
We can see that the algorithm converges linearly with complexity ${\cal O}(T_{1}+T_{2})={\cal O}(\log(\epsilon^{-1}))$. In this way, obtaining
@@
requires complexity $T={\cal O}(\log((\sqrt{\epsilon})^{-1}))={\cal O}(\log(\epsilon^{-1}))$. This completes the proof. ∎
@@ E.1 Problem formulation
$\min_{c}~{}~{}~{}\mathcal{L}_{\mathcal{D}_{\rm val}}(w^{*},b^{*})=$ $\sum_{(z_{{\rm val},},l_{{\rm val},})\in\mathcal{D}_{\rm val}}\exp\left({1-l_{{\rm val},}\left(z_{{\rm val},}^{\top}w^{*}+b^{*}\right)}\right)+\frac{1}{2}\|c\|^{2}$ (63a)
@@
The first term of upper-level objective (63a) is a validation loss, evaluated on the validation set $\mathcal{D}_{\rm val}$, the second is a regulation term on the upper-level variable $c$; the lower-level problem is designed to train the parameters of the hyperplane on the training set $\mathcal{D}_{\rm tr}$, with the soft margin violation $\xi_{i}$ being upper bounded by the hyperparameter $c_{i}$ [cf. (63d), which are the CCs]. The lower-level objective focuses on maximizing the margin by minimizing $\|w\|^{2}$ while allowing for violations $\xi$ to the separating hyperplane, which are regulated by the hyperparameter $c$. The BLO formulation aims to adjust the hyperparameter $c$ using the validation loss (upper-level objective), ensuring that the model parameters (lower-level variables) are optimal for the training dataset.
@@ E.2 Experiment details
In this section, we present detailed experimental results for the training of the SVM model in (63) using our BLOCC algorithm. We compare our algorithm to two baselines, LV-HBA [72] and GAM [69], both designed for BLO problems with inequality CCs. The GAM algorithm often encounters issues with matrix inversion, as the matrix required for the GAM algorithm to solve the problem in (63) can be singular, preventing it from finding a solution. Consequently, only BLOCC and LV-HBA solve the problem in (63), while GAM uses a different formulation as introduced in [69]. We compare the algorithms based on validation loss (the first term in the upper-level objective) and classification accuracy. These metrics are computed for both validation and test datasets. We evaluate the algorithms on two datasets: diabetes [20] and fourclass [29].
@@
(c) Lower level objective.
@@
(c) Lower level objective.
@@ F.1 Problem formulation
In transportation network planning, the planning operator is to construct a new transportation network (upper level) connecting a collection of stations $\mathcal{S}$, and the passengers will decide whether to use the new network (lower level), considering the options given by the new network and existing alternatives (constraints). The goal is to design a network that maximizes the operator’s benefit, knowing the passengers will make rational choices depending on the given design.
@@ 3. A3)
  • $\mathcal{Y}=[0,1]^{|\mathcal{K}|}\times[0,1]^{|\mathcal{A}||\mathcal{K}|}$ represents the domain of $y$.
@@
$\min_{x\in\mathcal{X}}f(x,y^{*}_{g}(x)):=-\Bigg{(}\sum_{\forall(o,d)\in\mathcal{K}}m^{od}y^{\rm od*}(x)-\sum_{\forall(i,j)\in\mathcal{A}}c_{ij}x_{ij}\Bigg{)},$ (64)
@@
For the lower level, we model the passenger’s behavior by finding the flow variables that maximize utility and minimizes flow entropy cost.
@@
$\qquad+\underbrace{\sum_{(o,d)\in\mathcal{K}}w^{od}y^{od}(\log(y^{od})-1)+\sum_{(o,d)\in\mathcal{K}}w^{od}(1-y^{od})(\log(1-y^{od})-1)}_{\rm flow~{}entropy~{}cost}\Bigg{)}.$
The passengers’ utility considers the time cost of choosing the new network and the existing network. The users set the flow variables $y$ so that the transportation network yielding a higher utility is preferred. Here, the probability of chosing new network is modeled by a logistic (softmax) model. However, setting the objective as a simple linear utility maximization would lead to an all-or-nothing policy, which is not the behavior observed in practice. Hence, the second term is brought to consider. The approach considered here is to formulate an objective given by the Legendre transform (cf. Definition 5) of the softmax sharing (see [53, 56, 55] for additional details). Intuitively, this means that rather than imposing the softmax sharing a fortiori, we formulate a convex problem whose KKT conditions lead to the softmax sharing. Using the fact that the Legendre transform of an exponential $e^{y}$ is the negative entropy function $y(\log(y)-1)$, the lower-level objective is traditional as in [56].
@@ 3. S3)
    The BLOCC algorithm produces two lower level variables: $y_{F,T}^{T_{F}}$ and $y_{g,T}^{T_{g}}$. While Theorems 1 and 2 guarantee that $(x_{T},y_{F,T}^{T_{F}})$ is an $\epsilon$-approximate solution, the pair $(x_{T},y_{g,T}^{T_{g}})$ is strictly feasible, meaning that $y_{g,T}^{T_{g}}=y_{g}^{*}(x_{T})$. Since strict feasibility is important in the transportation network, this section will report the results using both $y_{F,T}^{T_{F}}$ and $y_{g,T}^{T_{g}}$.
Figure 7: Negative upper-level objective $-f(x_{t},y_{t})$ for a 3-node network design problem; Solid lines show mean value of $-f(x_{t},y_{g,t}^{T_{g}})$ with the shaded region as a standard deviation; Dashed lines show the mean value of $-f(x_{t},y_{F,t}^{T_{F}})$ with the shaded region as a standard deviation; Three different $\gamma$ values (red, purple, blue) and fixed stepsize $\eta=1.6e-4$; The orange color represents the result of the LV-HBA algorithm.
Figure 8: Optimality gap of the lower-level problem for a 3-node network design problem $g(x_{t},y_{t})-g(x_{t},y^{*}_{t})$. Solid lines represent the mean value of the 10 realizations of the upper-level variables, dashed lines represent $g(x_{t},y_{F,t}^{T_{F}})-g(x_{t},y^{*}_{t})$, and the shaded region is the standard deviation. Three different $\gamma$ values are represented in our algorithm, and fixed stepsize $\eta=1.6e-4$.
@@ F.2 Numerical results for the 3-node network
For BLOCC, we set the stepsize to $\eta=1.6\times 10^{-4}$ and analyze the algorithm’s convergence for three different values of $\gamma$: $\gamma=2$, $\gamma=3$, and $\gamma=4$. The upper-level objective values are computed using $-f(x_{t},y_{g,t}^{T_{g}})$ and $-f(x_{t},y_{F,t}^{T_{F}})$. The results are presented in Figures 8 and 8.
Figure 8 illustrates the performance of our BLOCC algorithm over time. The orange line represents the evolution of $-f(x_{t},y_{t})$ for the LV-HBA algorithm, while the other six lines represent different implementations of BLOCC. Each color represents a different value of $\gamma$; solid lines represent $-f(x_{t},y_{g,t}^{T_{g}})$ and dashed lines represent $-f(x_{t},y_{F,t}^{T_{F}})$. Each simulation is conducted 10 times with 10 different random initializations of the upper-level variables, and both the mean and the standard deviation values are displayed. The results indicate that all versions of our BLOCC algorithm converge in less than 10 seconds, while LV-HBA shows slight fluctuations even after running for more than 50 seconds. This may be attributed to the fact that the LV-HBA algorithm requires a joint projection into $\{\mathcal{X}\times\mathcal{Y}:g^{c}(x,y)\leq 0\}$ at each iteration, involving 42 variables and 6 CCs.
@@
Figure 9: Negative upper-level objective $-f(x_{t},y_{t})$ for a 9-node network design problem; Solid lines show the mean value of $-f(x_{t},y_{g,t}^{T_{g}})$ with the shaded region as a standard deviation; Dashed lines show the mean value of $-f(x_{t},y_{F,t}^{T_{F}})$ with the shaded region as a standard deviation; Three different $\gamma$ values (red, purple, blue) and fixed stepsize $\eta=1.6e-4$; The orange color represents the result of the LV-HBA algorithm.
Figure 10: Negative upper-level objective $-f(x_{t},y_{t})$ for a metro network design problem in Seville, Spain, for 2 random initializations of the upper-level variables. Solid lines represent the mean of $-f(x_{t},y_{g,t}^{T_{g}})$, and the shaded region is the standard deviation. The dashed lines represent the mean of $-f(x_{t},y_{F,t}^{T_{F}})$, and the shaded region is the standard deviation. Three different $\gamma$ values are tested with a fixed stepsize $\eta=1.6e-4$.

### v2

@@ 4.1 Toy example
Table 2: Numerical results on the training outcome of our BLOCC in comparison with LV-HBA [75] and GAM [72]. The first row represents accuracy mean $\pm$ standard deviation, and the second row between brackets represents the running time until the upper-level objective’s update is smaller than 1e-5.
@@ 4.2 Hyperparameter optimization for SVM
We test the performance of our algorithm BLOCC with $\gamma=12$ when training a linear SVM model on the diabetes [21] and fourclass datasets [30]. The model is trained via the BLO formulation (1) with inequality CCs; see the details in Appendix E. We compared performance with two baselines, LV-HBA [75] and GAM [72].
@@ 4.3 Transportation network design problem
BLO is particularly relevant in transportation, where network planning must consider different time horizons and actors. Those problems are large-scale with a large number of upper- and lower-level variables and CCs, challenging the use of traditional BLO techniques. As a result, our final experiment considers a network design problem, where we act as an operator (whose utility is modeled as the upper-level objective) that considers the passengers’ behavior, modeled in the lower-level. We considered three networks: two synthetic networks of 3 and 9 nodes, respectively, and a real-world network of 26 nodes in the city of Seville, Spain. Further details about the formulation and the experiment can be found in Appendix F.
In this experiment, we only compare our BLOCC with LV-HBA, as GAM and BVFSM relies on the hypergradient of lower-level and can not tackle the problem with lower-level domain constraint $\mathcal{Y}$. Table 3 summarizes the experiment results. We can see that LV-HBA failed to work efficiently, especially for large networks. This is mainly because the increased constraints render the projection step impracticable. Our BLOCC, in contrast, is much faster, and it successfully converges even with large real-world networks. We provided computational complexity analysis in Appendix H and we can conclude that our BLOCC is robust to large-scale problems.
@@ B.1 Proof of Lemma 1
As $g(x,y)$ is $\alpha_{g}$-strongly convex in $y$ and $g^{c}(x,y)$ is convex in $y$, we know $g(x,y)+\langle\mu_{g}^{*}(x),g^{c}(x,y)\rangle$ is $\alpha_{g}$-strongly convex in $y$, where the modulus $\alpha_{g}$ is independent of $x$. Therefore, according to Appendix F and G in [38], and Theorem 3.3 in [19], the quadratic growth in statement 1 can be concluded.
@@ D.2 Proof of Theorem 3
Remark 2.
Under the same assumptions as in Theorem 3, we cam choose $\eta_{g,1}\leq(l_{g,1}+l_{g^{c},1})^{-1}$, $\eta_{g,2}\leq\frac{\alpha_{g}}{l_{g^{c},0}}$ as stepsizes for running the accelerated version of Algorithm 2 to solve (7), and $\eta_{F,1}\leq(l_{f,1}+\gamma l_{g,1}+l_{g^{c},1})^{-1}$, $\eta_{F,2}\leq\frac{\gamma\alpha_{g}-l_{f,1}}{l_{g^{c},0}}$ as the ones for (5).
@@ D.3 Proof of Theorem 4
with complexity $T={\cal O}(\ln(\epsilon^{-1}))$. Here, $(\mu^{*},y^{*})=\arg\max_{\mu\in\mathbb{R}^{d_{c}}}\min_{y\in\mathbb{R}^{d_{y}}}L(\mu,y)$.
@@
Notice ${\cal O}(\alpha_{h_{2}}^{-1})<{\cal O}(1)$ and $P_{0}={\cal O}(\epsilon^{-1})$ as $\rho=\epsilon^{-1}$. In this way, there exist $T_{1}={\cal O}(\ln(\epsilon^{-1}))$ such that for all $t>T_{1}$, $(1-c)^{t}P_{0}={\cal O}(1)$ and ${\cal O}(\alpha_{h_{2}}^{-1})\leq{\cal O}(1)$. Accordingly, we can achieve
@@
This is an iteration outcome using (50) in Lemma 12, (62), and the fact that $\eta_{1}\alpha_{h_{2}}/2={\cal O}(\frac{\alpha_{h_{2}}}{l_{h_{2},1}})={\cal O}(1)$. In this way, for another $T_{2}={\cal O}(\ln(\epsilon^{-1}))$ steps, we have
@@
We can see that the algorithm converges linearly with complexity ${\cal O}(T_{1}+T_{2})={\cal O}(\ln(\epsilon^{-1}))$. In this way, obtaining
@@
requires complexity $T={\cal O}(\ln((\sqrt{\epsilon})^{-1}))={\cal O}(\ln(\epsilon^{-1}))$. This completes the proof. ∎
Remark 3.
Under the same assumptions as in Theorem 4, we cam choose $\eta_{g,2}\lesssim(l_{g,1})^{-1}$, $\eta_{g,1}\leq\alpha_{g}^{2}\epsilon_{g}(2s_{\max}\alpha_{g}+s_{\max}^{2}\epsilon^{-1})^{-1}\eta_{g,2}$ as stepsizes for running the single-loop version of Algorithm 2 to solve (7), and $\eta_{F,2}\lesssim(l_{f,1}+\gamma l_{g,1})^{-1}$, $\eta_{F,2}\leq(\gamma\alpha_{g}-l_{f,1})^{2}\epsilon_{F}(2s_{\max}(\gamma\alpha_{g}-l_{f,1})+s_{\max}^{2}\epsilon^{-1})^{-1}\eta_{F,2}$ as the ones for (5).
@@ E.1 Problem formulation
$\min_{c}~{}~{}~{}\mathcal{L}_{\mathcal{D}_{\rm val}}(w^{*},b^{*})=$ $\sum_{(z_{{\rm val}},l_{{\rm val}})\in\mathcal{D}_{\rm val}}\exp\left({1-l_{{\rm val}}\left(z_{{\rm val}}^{\top}w^{*}+b^{*}\right)}\right)+\frac{1}{2}\|c\|^{2}$ (63a)
@@
The first term of upper-level objective (63a) is a validation loss, evaluated on the validation set $\mathcal{D}_{\rm val}$, the second is a regulation term on the upper-level variable $c$; the lower-level problem is designed to train the parameters of the hyperplane on the training set $\mathcal{D}_{\rm tr}$, with the soft margin violation $\xi_{i}$ being upper bounded by the hyperparameter $c_{i}$ [cf. (63d), which are the CCs]. The lower-level objective (63b) focuses on maximizing the margin by minimizing $\|w\|^{2}$ while allowing for violations $\xi$ to the separating hyperplane, which are regulated by the hyperparameter $c$. The BLO formulation aims to adjust the hyperparameter $c$ using the validation loss (upper-level objective), ensuring that the model parameters (lower-level variables) are optimal for the training dataset.
@@ E.2 Experiment details
In this section, we present detailed experimental results for the training of the SVM model in (63) using our BLOCC algorithm. We compare our algorithm to two baselines, LV-HBA [75] and GAM [72], both designed for BLO problems with inequality CCs. We use $\gamma=12$ and $\eta=0.01$ for running our BLOCC and we apply $\alpha=0.01$, $\gamma_{1}=0.1$, $\gamma_{2}=0.1$, $\eta=0.001$ for LV-HBA and $\alpha=0.05$, $\epsilon=0.005$ for GAM respectively. The hyper-parameters for LV-HBA and GAM are the ones used in the SVM experiments in their paper. The GAM algorithm often encounters issues with matrix inversion, as the matrix required for the GAM algorithm to solve the problem in (63) can be singular, preventing it from finding a solution. Consequently, only BLOCC and LV-HBA solve the problem in (63), while GAM uses a different formulation as introduced in [72]. We compare the algorithms based on validation loss (the first term in the upper-level objective) and classification accuracy. These metrics are computed for both validation and test datasets. We evaluate the algorithms on two datasets: diabetes [21] and fourclass [30].
@@
(c) Lower-level objective.
@@
(c) Lower-level objective.
@@ F.1 Problem formulation
In transportation network planning, the planning operator is to construct a new transportation network (upper level) connecting a collection of stations $\mathcal{S}$, and the passengers will decide whether to use the new network (lower-level), considering the options given by the new network and existing alternatives (constraints). The goal is to design a network that maximizes the operator’s benefit, knowing the passengers will make rational choices depending on the given design.
@@ 3. A3)
  • $\mathcal{Y}=[\varepsilon,1-\varepsilon]^{|\mathcal{K}|}\times[\varepsilon,1-\varepsilon]^{|\mathcal{A}||\mathcal{K}|}$, where $\varepsilon$ is a small positive number set by the designer of the network, represents the domain of $y$.
@@
$\min_{x\in\mathcal{X}}f(x,y^{*}_{g}(x)):=-\Bigg{(}\underbrace{\sum_{\forall(o,d)\in\mathcal{K}}m^{od}y^{\rm od*}(x)}_{\text{profit}}-\underbrace{\sum_{\forall(i,j)\in\mathcal{A}}c_{ij}x_{ij}}_{cost}\Bigg{)},$ (64)
@@
For the lower-level, we model the passenger’s behavior by finding the flow variables that maximize utility and minimizes flow entropy cost.
@@
$\qquad+\underbrace{\sum_{(o,d)\in\mathcal{K}}w^{od}y^{od}(\ln(y^{od})-1)+\sum_{(o,d)\in\mathcal{K}}w^{od}(1-y^{od})(\ln(1-y^{od})-1)}_{\rm flow~{}entropy~{}cost}\Bigg{)}.$
The passengers’ utility considers the time cost of choosing the new network and the existing network. The users set the flow variables $y$ so that the transportation network yielding a higher utility is preferred. Here, the probability of chosing new network is modeled by a logistic (softmax) model. However, setting the objective as a simple linear utility maximization would lead to an all-or-nothing policy, which is not the behavior observed in practice. Hence, the second term is brought to consider. The approach considered here is to formulate an objective given by the Legendre transform (cf. Definition 5) of the softmax sharing (see [55, 58, 57] for additional details). Intuitively, this means that rather than imposing the softmax sharing a fortiori, we formulate a convex problem whose KKT conditions lead to the softmax sharing. Using the fact that the Legendre transform of an exponential $e^{y}$ is the negative entropy function $y(\ln(y)-1)$, the lower-level objective is traditional as in [58].
@@ 3. S3)
    The BLOCC algorithm produces two lower-level variables: $y_{F,T}^{T_{F}}$ and $y_{g,T}^{T_{g}}$. While Theorems 1 and 2 guarantee that $(x_{T},y_{F,T}^{T_{F}})$ is an $\epsilon$-approximate solution, the pair $(x_{T},y_{g,T}^{T_{g}})$ is strictly feasible, meaning that $y_{g,T}^{T_{g}}=y_{g}^{*}(x_{T})$. Since strict feasibility is important in the transportation network, this section will report the results using both $y_{F,T}^{T_{F}}$ and $y_{g,T}^{T_{g}}$.
Figure 7: Upper-level objective $f(x_{t},y_{t})$ for a 3-node network design problem; Solid lines show mean value of $f(x_{t},y_{g,t}^{T_{g}})$ with the shaded region as a standard deviation; Dashed lines show the mean value of $f(x_{t},y_{F,t}^{T_{F}})$ with the shaded region as a standard deviation; Three different $\gamma$ values (red, purple, blue) and fixed stepsize $\eta=1.6\times 10^{-4}$; The orange color represents the result of the LV-HBA algorithm.
Figure 8: Optimality gap of the lower-level problem for a 3-node network design problem $g(x_{t},y_{t})-g(x_{t},y^{*}_{t})$. Solid lines represent the mean value of the 10 realizations of the upper-level variables, dashed lines represent $g(x_{t},y_{F,t}^{T_{F}})-g(x_{t},y^{*}_{t})$, and the shaded region is the standard deviation. Three different $\gamma$ values are represented in our algorithm, and fixed stepsize $\eta=1.6\times 10^{-4}$.
@@ F.2 Numerical results for the 3-node network
For BLOCC, we set the stepsize to $\eta=1.6\times 10^{-4}$ and analyze the algorithm’s convergence for three different values of $\gamma$: $\gamma=2$, $\gamma=3$, and $\gamma=4$. The upper-level objective values are computed using $f(x_{t},y_{g,t}^{T_{g}})$ and $f(x_{t},y_{F,t}^{T_{F}})$. The results are presented in Figures 8 and 8.
Figure 8 illustrates the performance of our BLOCC algorithm over time. The orange line represents the evolution of $f(x_{t},y_{t})$ for the LV-HBA algorithm, while the other six lines represent different implementations of BLOCC. Each color represents a different value of $\gamma$; solid lines represent $f(x_{t},y_{g,t}^{T_{g}})$ and dashed lines represent $f(x_{t},y_{F,t}^{T_{F}})$. Each simulation is conducted 10 times with 10 different random initializations of the upper-level variables, and both the mean and the standard deviation values are displayed. The results indicate that all versions of our BLOCC algorithm converge in less than 10 seconds, while LV-HBA shows slight fluctuations even after running for more than 50 seconds. This may be attributed to the fact that the LV-HBA algorithm requires a joint projection into $\{\mathcal{X}\times\mathcal{Y}:g^{c}(x,y)\leq 0\}$ at each iteration, involving 42 variables and 6 CCs.
@@
Figure 9: Upper-level objective $f(x_{t},y_{t})$ for a 9-node network design problem; Solid lines show the mean value of $f(x_{t},y_{g,t}^{T_{g}})$ with the shaded region as a standard deviation; Dashed lines show the mean value of $f(x_{t},y_{F,t}^{T_{F}})$ with the shaded region as a standard deviation; Three different $\gamma$ values (red, purple, blue) and fixed stepsize $\eta=1.6\times 10^{-4}$; The orange color represents the result of the LV-HBA algorithm.
Figure 10: Upper-level objective $f(x_{t},y_{t})$ for a metro network design problem in Seville, Spain, for 2 random initializations of the upper-level variables. Solid lines represent the mean of $f(x_{t},y_{g,t}^{T_{g}})$, and the shaded region is the standard deviation. The dashed lines represent the mean of $f(x_{t},y_{F,t}^{T_{F}})$, and the shaded region is the standard deviation. Three different $\gamma$ values are tested with a fixed stepsize $\eta=1.6\times 10^{-4}$.